Does the quadratic function y = -12(x+2)^2-7 have it's vertex at (-2,7) and open down? If not,write a function that meets the criteria.
Answer by venugopalramana(3286)

Does the quadratic function y = -12(x+2)^2-7
have it's vertex at (-2,7)
X=-2 VALUE IS OK.BUT THEN WE GET Y=-7...SO VERTEX IS (-2,-7)
SINCE Y=-7-12(X+2)^2...HAS ITS MAXIMUM WHEN X=-2,VERTEX REPRESENTS A MAXIMUM...THAT IS OPEN DOWN
and open down?...OK
If not,write a function that meets the criteria.
Y=7-12(X+2)^2...HAS VERTEX AT (-2,7) AND OPENS DOWN WARDS.